# Hartman Effect from a Geometrodynamic Extension of Bohmian Mechanics

### v2

@@ 6 Momentum, Position, and Tunneling Time
But what is the physical or mathematical reason behind this condition? The answer to this question arises when analyzing the expression $\alpha_{1}=1/\cosh{(2\sigma R)}=\mbox{sech}(2\sigma R)$ introduced in Section 4, from which we derive $\sigma=\mbox{arsech}(\alpha_{1})/2R$ and consequently $\mbox{arsech}(\alpha_{1})\ll 2R$. It is worthwhile mentioning that as the quantity $\mbox{arsech}(\alpha_{1})$ is a dimensionless constant, the above expression does not consider the units of $R$ or, in the particular case, $R_{e}$; that is, we only aim to establish a comparison in orders of magnitude.
Therefore, in order to provide a physical meaning to this last mathematical condition, particles of small dimensions, such as an electron, are considered, whose radius has the value $R_{e}=2.8178402894(58)\times 10^{-13}\,\text{m}$. Comparing it with the Bohr radius ($a_{0}=5.29177210903(86)\times 10^{-11}\,\text{m}$), we obtain the relation $a_{0}=188R_{e}$. Thus, it is sufficient to choose an $\alpha_{1}$ such that $R_{e}<\mbox{arsech}(\alpha_{1})\ll 376R_{e}$ to obtain an appropriate parameter $\sigma$. Therefore, by solving the differential equations (56), (57), and (58) (see the Appendix), we obtain the expressions
@@
Where $\varrho_{I}$, $\varrho_{II}$, $\varrho_{III}$ are arbitrary constants, $\vartheta_{3}^{\prime}=\mathcal{B}\vartheta_{3}/\mathcal{A}$, $\varsigma_{I}^{0}=\vartheta_{3}^{\prime}\vartheta_{4}[\vartheta_{3}^{\prime}-\vartheta_{1}]/\{4[\vartheta_{3}^{\prime}\vartheta_{4}-\vartheta_{1}\vartheta_{2}]\}$, $\varsigma_{I}^{1}=\vartheta_{3}^{\prime}\vartheta_{1}[\vartheta_{3}^{\prime}-\vartheta_{1}][\vartheta_{2}-\vartheta_{4}][\vartheta_{3}^{\prime}(1-\vartheta_{4})-\vartheta_{1}(1+\vartheta_{2})]^{\frac{1}{2}}/\{2[\vartheta_{3}^{\prime}\vartheta_{4}-\vartheta_{1}\vartheta_{2}][\vartheta_{3}^{\prime}(1-\vartheta_{4})-\vartheta_{1}(1-\vartheta_{2})]^{\frac{3}{2}}\}$, $\varsigma_{I}^{2}=\{[\vartheta_{3}^{\prime}(1-\vartheta_{4})-\vartheta_{1}(1+\vartheta_{2})]/[\vartheta_{3}^{\prime}(1-\vartheta_{4})-\vartheta_{1}(1-\vartheta_{2})]\}^{\frac{1}{2}}$, $\varsigma_{II}^{0}=-[1+\iota_{1}^{\prime}+\iota_{2}^{\prime}+\iota_{3}^{\prime}+\iota_{4}^{\prime}+\iota_{5}^{\prime}+\iota_{6}^{\prime}]/[\iota_{1}^{\prime\prime}+\iota_{2}^{\prime\prime}+\iota_{3}^{\prime\prime}+\iota_{4}^{\prime\prime}+\iota_{5}^{\prime\prime}-\iota_{6}^{\prime\prime}-1]$, $\varsigma_{II}^{1}=-[-(1+\iota_{1}^{\prime}+\iota_{2}^{\prime}+\iota_{3}^{\prime}+\iota_{4}^{\prime}+\iota_{5}^{\prime}+\iota_{6}^{\prime})(\iota_{1}^{\prime\prime}+2\iota_{2}^{\prime\prime}+3\iota_{3}^{\prime\prime}+4\iota_{4}^{\prime\prime}+5\iota_{5}^{\prime\prime}-6\iota_{6}^{\prime\prime})+(\iota_{1}^{\prime\prime}+\iota_{2}^{\prime\prime}+\iota_{3}^{\prime\prime}+\iota_{4}^{\prime\prime}+\iota_{5}^{\prime\prime}-\iota_{6}^{\prime\prime}-1)(\iota_{1}^{\prime}+2\iota_{2}^{\prime}+3\iota_{3}^{\prime}+4\iota_{4}^{\prime}+5\iota_{5}^{\prime}+6\iota_{6}^{\prime})]/[\iota_{1}^{\prime\prime}+\iota_{2}^{\prime\prime}+\iota_{3}^{\prime\prime}+\iota_{4}^{\prime\prime}+\iota_{5}^{\prime\prime}-\iota_{6}^{\prime\prime}-1]^{2}$, $\iota_{1}^{\prime\prime}=\iota_{0}\iota_{1}-\iota_{1}^{\prime}$, $\iota_{2}^{\prime\prime}=\iota_{0}\iota_{2}-\iota_{2}^{\prime}$, $\iota_{3}^{\prime\prime}=\iota_{0}\iota_{3}-\iota_{3}^{\prime}$, $\iota_{4}^{\prime\prime}=\iota_{0}\iota_{4}-\iota_{4}^{\prime}$, $\iota_{5}^{\prime\prime}=\iota_{0}\iota_{5}-\iota_{5}^{\prime}$, $\iota_{6}^{\prime\prime}=\iota_{6}^{\prime}$, $\varsigma_{III}^{0}=\vartheta_{7}^{\prime}\vartheta_{8}[\vartheta_{7}^{\prime}-\vartheta_{5}]/\{4[\vartheta_{7}^{\prime}\vartheta_{8}-\vartheta_{5}\vartheta_{6}]\}$, $\varsigma_{III}^{1}=\vartheta_{7}^{\prime}\vartheta_{5}[\vartheta_{7}^{\prime}-\vartheta_{5}][\vartheta_{6}-\vartheta_{8}][\vartheta_{7}^{\prime}(1-\vartheta_{8})-\vartheta_{5}(1+\vartheta_{6})]^{\frac{1}{2}}/\{2[\vartheta_{7}^{\prime}\vartheta_{8}-\vartheta_{5}\vartheta_{6}][\vartheta_{7}^{\prime}(1-\vartheta_{8})-\vartheta_{5}(1-\vartheta_{6})]^{\frac{3}{2}}\}$ and $\varsigma_{III}^{2}=\{[\vartheta_{7}^{\prime}(1-\vartheta_{8})-\vartheta_{5}(1+\vartheta_{6})]/[\vartheta_{7}^{\prime}(1-\vartheta_{8})-\vartheta_{5}(1-\vartheta_{6})]\}^{\frac{1}{2}}$. Then, as the expressions presented in equations (59)-(61) do not allow for solving one variable in terms of the other to obtain the general equation of trajectories for all time $t$, it is possible, at least, to provide an idea of such trajectories for the stationary case by plotting a set of isochronous curves. These curves are obtained by recalling that $\vartheta_{3}^{\prime}=\mathcal{B}\vartheta_{3}/\mathcal{A}=[A-B]\vartheta_{3}/[A+B]$, where $A$ and $B$ are arbitrary constants. This allows rewriting Eq. (59) as the sum of two terms, one positive corresponding to incident particles and one negative corresponding to reflected particles. Thus, the graphical representation of the expressions (59)-(61) is obtained, as shown in Fig. 2.
@@
Figure 2: Schematic representation of the trajectories followed by particles incident from the left on the potential barrier.
In Figure 2, trajectories of incident and reflected particles are depicted in blue and yellow, respectively, in region $I$. In region $II$, trajectories of tunneling particles are represented by a blue-dashed line. In region $III$, trajectories of transmitted particles are shown. It is essential to highlight that the trajectories presented in this plot correspond to the curves of the incident functions $f^{(i)}(x,t)=0.25(2.41x-nt)+\tan^{-1}{\{10\tan{[(2.41x-nt)/2]}\}}/10+\varrho_{I}^{(i)}$, reflected $f^{(r)}(x,t)=-0.25(2.41x-nt)-\tan^{-1}{\{10\tan{[(2.41x-nt)/2]}\}}/10+\varrho_{I}^{(r)}$, tunneling $f^{(tu)}(x,t)=0.0995(x-nt)+\varrho_{II}^{(tu)}$, and transmitted $f^{(tr)}(x,t)=0.25[2.41x-nt]+\tan^{-1}{\{10\tan{[(2.41x-nt)/2]}\}}/10+\varrho_{III}^{(tr)}$ (with arbitrary chosen values for the constants in each expression). Taking $n=2$, $t=1$, and the sets of values $\varrho_{I}^{(i)}=\{4.65,5.65,6.65,7.65,8.65,9.65\}$, $\varrho_{I}^{(r)}=\{-1.75,-0.75,0.25,1.25\}$, $\varrho_{II}^{(tu)}=\{1.75,2.75,3.75,4.75\}$, $\varrho_{III}^{(tr)}=\{1.3,2.3,3.3,4.3\}$, where the closest plots to the $x$-axis correspond to the first elements of each set of values and so on.
@@
whose covariant derivative [^3] Finally, following the line of clarifications made previously in sections 2, 3 and 4, we observe that in this work $()_{;\nu}=\partial_{\nu}()$ is used to denote the covariant derivative. allows us to obtain the expansion of the volume elements [38] produced by the bubble
@@
and whose component $u^{1}=v$ mathematically expresses the speed with which the ‘warp’ bubble moves in the $ox$ direction. Consequently, due to the previous consideration $\alpha_{1}=2\rightarrow\sigma\ll 1$, we have that
$v=-v_{s}\left[-\frac{v_{s}}{v_{s}(1+2)}\right]=\frac{v_{s}}{3}.$ (64)
@@
This leads to the following two implications. First, the fact that $ds^{2}<0$ implies that the curve representing the trajectory followed by the particle is of a time-like nature. Second, by rewriting Eq. (66), considering the definition of proper time, $d\tau^{2}=-ds^{2}$, we obtain
@@
which implies that the metric proposed in Refs. 38, 48 does not generate temporal dilations between the time measured with respect to the particle and the coordinate time.
For this reason, we consider a particle moving inside the ‘warp’ bubble [47] following the trajectory $x=x^{\prime}+vt=x^{\prime}+nt/3$ (with $v_{s}=n\rightarrow v=n/3$) and again $\sigma\ll 1$. With this, we can rewrite Eq. (60) as follows
$\varrho_{II}=\varsigma_{II}^{0}\biggl{(}x^{\prime}-\frac{2n}{3}t\biggr{)}.$ (68)
Then, by evaluating Eq. (68) at the points $\mathcal{P}_{0}(-a/2,t_{0})$ and $\mathcal{P}_{1}(a/2,t_{1})$, we get
$\varrho_{II}=\varsigma_{II}^{0}\biggl{(}-\frac{a}{2}-\frac{2n}{3}t_{0}\biggr{)}$ (69)
@@
$\varrho_{II}=\varsigma_{II}^{0}\biggl{(}\frac{a}{2}-\frac{2n}{3}t_{1}\biggr{)},$ (70)
which, when subtracted, gives us
$\Delta t=\frac{3}{2n}a.$ (71)
@@
In other words, the tunneling time (71), graphically represented in Fig. 3, depends on the barrier width, as expected. That is, with an increase in the potential barrier width, the tunneling time could increase depending on the value of the tunneling velocity. This velocity could even reach superluminal regimes due to the distortion of spacetime, as seen in Eq. (11) or in Eq. (22), generated by the quantum potential of Eq. (51).
As a result, by choosing $v_{s}=n$ in this model, there is no saturation in the tunneling times obtained in Refs. 1, 13, 14, 19. Therefore, it can be said that the solution Eq. (48), constructed in this work, eliminates the apparent paradox by considering the Alcubierre-type spacetime inside the potential barrier, where particles follow geodesic trajectories. However, despite the general expression
$\varrho_{II}=\varsigma_{II}^{0}(v_{s})(x-v_{s}t),$ (72)
could lead to expressions for tunneling times that exhibit saturation (as a consequence of the freedom in choosing $v_{s}$), the solution (48) still reconciles the apparent contradiction between quantum mechanics and general relativity by considering a spacetime that allows for superluminal travel without violating the foundations of relativistic theory [38, 48, 35, 36]. Therefore, the Hartman effect [1] would be explained again as a distortion of spacetime generated by the quantum potential (51), which also explicitly depends on $v_{s}$.
Figure 3: Graphical representation of tunneling time as a linear function of the width of the potential barrier for different values of the parameter $n$.

### v3

@@ 6 Momentum, Position, and Tunneling Time
But what is the physical or mathematical reason behind this condition? The answer to this question arises when we understand that in this context we are trying to compare the value of $\sigma$ with respect to the radius of the warp bubble, that is, the dimension of the disturbance produced in space-time. However, since this disturbance is generated to transport particles at hyperluminal speeds, when thinking about an electron, the situation at first glance becomes complicated to interpret. Since in orthodox quantum theory the electron exhibits both corpuscular and wave-like behavior, described by the wave function, which does not express the physical size of the electron, but rather the probability of finding it in different regions of space.
Furthermore, the idea of “size” for an electron is limited by the probabilistic nature of quantum mechanics and by Heisenberg’s uncertainty principle. However, in some contexts, it is possible to speak of an “effective dimension” of the electron, related to the spatial extension of its wave function, for example: in an atom, the electron occupies an orbital with a certain shape and size. However, at the fundamental level of Bohmian mechanics, the electron is considered a point particle without internal structure (classical idea), whose dynamics are governed by the guiding wave. Therefore, to give a physical meaning to this last mathematical condition, we consider, without loss of generality, particles of small dimensions, such as an electron, whose classical radius has the value $R_{e}=2.8178402894(58)\times 10^{-13}\,\mbox{m}$. Comparing it with the Bohr radius ($a_{0}=5.29177210903(86)\times 10^{-11}\,\mbox{m}$), we obtain the relation $a_{0}=188R_{e}$.
In this sense, when analyzing the expression $\alpha_{1}=1/\cosh{(2\sigma R)}=\mbox{sech}(2\sigma R)$ introduced in Section 4, from which we derive $\sigma=\mbox{arsech}(\alpha_{1})/2R$ and consequently $\mbox{arsech}(\alpha_{1})\ll 2R$. It is worth mentioning that, since the quantity $\mbox{arsech}(\alpha_{1})$ is a dimensionless constant, the expression above does not consider the units of $R$ or, in this particular case, $R_{e}$, that is, we only aim to establish a comparison in orders of magnitude. Thus, it is enough to choose a $\alpha_{1}$ such that $R_{e}<\mbox{arsech}(\alpha_{1})\ll 376R_{e}$ to obtain an appropriate parameter $\sigma$. Therefore, by solving the differential equations (55), (56), and (57) (see the Appendix), we obtain the expressions
@@
Where $\varrho_{I}$, $\varrho_{II}$, $\varrho_{III}$ are arbitrary constants, the coefficients in equations (58)-(60) are defined as follows:
$\varsigma_{I}^{0}$ $=$ $\frac{\vartheta_{10}[1-\vartheta_{11}]}{1-\vartheta_{12}},$
$\varsigma_{I}^{1}$ $=$ $\frac{2\vartheta_{10}[\vartheta_{11}-\vartheta_{12}]}{\sqrt{\vartheta_{12}}[1-\vartheta_{12}]},$
$\varsigma_{I}^{2}$ $=$ $\frac{1}{\sqrt{\vartheta_{12}}},$
$\varsigma_{III}^{0}$ $=$ $\frac{\vartheta_{10}^{\prime}[1-\vartheta_{11}^{\prime}]}{1-\vartheta_{12}^{\prime}},$
$\varsigma_{III}^{1}$ $=$ $\frac{2\vartheta_{10}^{\prime}[\vartheta_{11}^{\prime}-\vartheta_{12}^{\prime}]}{\sqrt{\vartheta_{12}^{\prime}}[1-\vartheta_{12}^{\prime}]},$
$\varsigma_{III}^{2}$ $=$ $\frac{1}{\sqrt{\vartheta_{12}^{\prime}}},$
$\vartheta_{3}^{\prime}$ $=$ $\frac{\mathcal{B}\vartheta_{3}}{\mathcal{A}},$
$\vartheta_{10}$ $=$ $\frac{\vartheta_{3}^{\prime}[1-\vartheta_{4}][\vartheta_{3}^{\prime}-\vartheta_{1}]}{\vartheta_{3}^{\prime}[1-\vartheta_{4}]-\vartheta_{1}[1-\vartheta_{2}]},$
$\vartheta_{11}$ $=$ $\frac{1+\vartheta_{4}}{1-\vartheta_{4}},$
$\vartheta_{12}$ $=$ $\frac{\vartheta_{3}^{\prime}[1+\vartheta_{4}]-\vartheta_{1}[1+\vartheta_{2}]}{\vartheta_{3}^{\prime}[1-\vartheta_{4}]-\vartheta_{1}[1-\vartheta_{2}]},$
$\vartheta_{10}^{\prime}$ $=$ $\frac{\vartheta_{7}[1-\vartheta_{8}][\vartheta_{7}-\vartheta_{5}]}{\vartheta_{7}[1-\vartheta_{8}]-\vartheta_{5}[1-\vartheta_{6}]},$
$\vartheta_{11}^{\prime}$ $=$ $\frac{1+\vartheta_{8}}{1-\vartheta_{8}},$
$\vartheta_{12}^{\prime}$ $=$ $\frac{\vartheta_{7}[1+\vartheta_{8}]-\vartheta_{5}[1+\vartheta_{6}]}{\vartheta_{7}[1-\vartheta_{8}]-\vartheta_{5}[1-\vartheta_{6}]},$
$\varsigma_{II}^{0}$ $=$ $-\frac{1+\sum_{j=1}^{6}\iota_{j}^{\prime}}{\sum_{j=1}^{6}\iota_{j}^{\prime\prime}-1},$
$\varsigma_{II}^{1}$ $=$ $-\frac{\mathcal{N}}{\mathcal{D}^{2}},$
$\mathcal{N}$ $=$ $(1+\sum_{j=1}^{6}\iota_{j}^{\prime})(6\iota_{6}^{\prime\prime}-\sum_{j=1}^{6}j\iota_{j}^{\prime\prime})+(\sum_{j=1}^{6}\iota_{j}^{\prime\prime}-1)(\sum_{j=1}^{6}j\iota_{j}^{\prime}),$
$\mathcal{D}$ $=$ $\sum_{j=1}^{6}\iota_{j}^{\prime\prime}-1,$
$\iota_{j}^{\prime\prime}$ $=$ $\iota_{0}\iota_{j}-\iota_{j}^{\prime}\;\;\mbox{(for $j=1,\dots,5$)},$
$\iota_{6}^{\prime\prime}$ $=$ $\iota_{6}^{\prime}.$
@@
Then, as the expressions presented in equations (58)-(60) do not allow for solving one variable in terms of the other to obtain the general equation of trajectories for all time $t$, it is possible, at least, to provide an idea of such trajectories for the stationary case by plotting a set of isochronous curves. These curves are obtained by recalling that $\vartheta_{3}^{\prime}=\mathcal{B}\vartheta_{3}/\mathcal{A}=[A-B]\vartheta_{3}/[A+B]$, where $A$ and $B$ are arbitrary constants. This allows rewriting Eq. (58) as the sum of two terms, one positive corresponding to incident particles and one negative corresponding to reflected particles. Thus, the graphical representation of the expressions (58)-(60) is obtained, as shown in Fig. 2.
@@
Figure 2: Graphical representation of the trajectories followed by particles incident from the left on the potential barrier.
In Fig. 2, the particle trajectories in the three regimes are presented: (i) region I (incident in blue and reflected in yellow), (ii) region II (tunneling, dashed blue line), and (iii) region III (transmitted). The curves correspond to the analytical solutions of Equations (57)-(59), which for the considered cases reduce to: $f^{(in)}(x,t)=(1.21x-1)/2+\tan^{-1}{[10\tan{(1.21x-1)}]}/10+\varrho_{I}^{(in)}$, $f^{(re)}(x,t)=-(1.21x-1)/2-\tan^{-1}{[10\tan{(1.21x-1)}]}/10+\varrho_{I}^{(re)}$, $f^{(tu)}(x,t)=0.0995(x-2)+\varrho_{II}^{(tu)}$, $f^{(tr)}(x,t)=(1.21x-1)/2+\tan^{-1}{[10\tan{(1.21x-1)}]}/10+\varrho_{III}^{(tr)}$.
We emphasize that the position $x_{s}(t)=2$, as well as the parameters $\mathcal{B}\Delta k/\mathcal{A}$ and $\Delta E$ in $\Xi_{I}$ and their analogues in $\Xi_{III}$, were chosen for convenience for the graphical representation. The constants $\varsigma_{j}^{\ell_{(\alpha)}}$ (with $\alpha\in\{in,re,tu,tr\}$, $\ell_{(\alpha)}\in\{0,1,2\}$ and $j\in\{I,II,III\}$) were determined by imposing continuity conditions at the potential barrier’s boundary points. The parameters $\varrho$ were discretized to represent different initial conditions: $\varrho_{I}^{(in)}\in\{4.65,5.65,\dots,9.65\}$, $\varrho_{I}^{(re)}\in\{-1.75,-0.75,0.25,1.25\}$, $\varrho_{II}^{(tu)}\in\{1.75,2.75,3.75,4.75\}$ and $\varrho_{III}^{(tr)}\in\{1.3,2.3,3.3,4.3\}$, with the trajectories closest to the $x$-axis corresponding to the smallest values in each set.
@@
whose covariant derivative allows us to obtain the expansion of the volume elements [45] produced by the bubble
@@
and whose component $u^{1}=v$ mathematically expresses the speed with which the ‘warp’ bubble moves in the $ox$ direction.
For this reason, when considering the low distortion condition ($\sigma\ll 1$) in the bubble profile $f(r_{s})$ (defined in Section 4), the warp bubble velocity $v$ in region $II$ (Eulerian observer velocity) is approximated by:
$v=-v_{s}f(r_{s})\approx-v_{s}\left.\left(\frac{\alpha_{0}}{v_{s}[1+\alpha_{1}\cosh{(2\sigma r_{s})}]}\right)\right|_{\sigma\ll 1}.$ (63)
Moreover, since $\sigma r_{s}\approx 0$ implies $\alpha_{1}=2$ and $\cosh(2\sigma r_{s})\approx 1$ from (63), we obtain:
$v\approx-\frac{\alpha_{0}}{1+\alpha_{1}}=-\frac{\alpha_{0}}{3}.$ (64)
Then, since $\alpha_{0}$ is defined by $\alpha_{0}=-[v_{s}\tanh{(2\sigma R)}]/[2\tanh{(\sigma R)}]$ (Section 4), and for $\sigma\ll 1$, $\tanh{(\sigma R)}\approx\sigma R$, we have:
$\alpha_{0}\approx-\frac{v_{s}(2\sigma R)}{2\sigma R}=-v_{s},$ (65)
which when substituted into (64) yields finally:
$v\approx-\frac{(-v_{s})}{3}=\frac{v_{s}}{3}.$ (66)
@@
This leads to the following implications. First, the fact that $ds^{2}<0$ implies that the curve representing the trajectory followed by the particle is of a time-like nature. Second, by rewriting Eq. (68), considering the definition of proper time, $d\tau^{2}=-ds^{2}$, we obtain
@@
which implies that the metric proposed in Refs. [45, 52] does not generate temporal dilations between the time measured with respect to the particle and the coordinate time.
For this reason, we consider a particle moving inside the “warp” bubble [51] following the trajectory $x=x^{\prime}+vt=x^{\prime}+v_{s}/3$ and again $\sigma\ll 1\rightarrow\sigma^{2}\approx 0$. Consequently, Eq. (59) can be rewritten as:
$\varrho_{II}=\varsigma_{II}^{0}\biggl(x^{\prime}-\frac{v_{s}}{3}t\biggr).$ (70)
Then, by evaluating Eq. (70) at the entry points $\mathcal{P}_{0}(-a/2,t_{0})$ and exit points $\mathcal{P}_{1}(a/2,t_{1})$, we obtain:
$\varrho_{II}=\varsigma_{II}^{0}\biggl(-\frac{a}{2}-\frac{v_{s}}{3}t_{0}\biggr)$ (71)
@@
$\varrho_{II}=\varsigma_{II}^{0}\biggl(\frac{a}{2}-\frac{v_{s}}{3}t_{1}\biggr),$ (72)
which, when subtracted, results in:
$\Delta t=\frac{3}{v_{s}}a.$ (73)
@@
This tunneling time depends on the barrier width, as expected; thus, an increase in the potential barrier width shall increase the tunneling time, depending on the value of the tunneling velocity.
In other words, a simple choice like $v_{s}=nc=n$ (since in natural units $c=1$) in (73) derived from the model implies that there is no apparent saturation in tunneling times as reported in the literature [1, 17, 18, 23]. However, it is essential to highlight that the solution (Equation 47) reproduces the Hartman effect through a different mechanism that must be carefully explained from the obtained results.
To achieve this, an asymptotic analysis of the function
$f(r_{s})=\frac{\alpha_{0}}{v_{s}[1+\alpha_{1}\cosh{(2\sigma r_{s})}]},$ (74)
must be performed, based on the hypothesis of small perturbations ($\sigma\ll 1$), where $r_{s}=x-x_{s}$, $\alpha_{0}\approx-v_{s}$ and $\alpha_{1}\approx 1$. Then, we firstly focus our attention on the inner region of the perturbation, $|r_{s}|<R$. For the regime of small perturbations, the function $f(r_{s})$ admits the following approximation
$f(r_{s})\approx\frac{\alpha_{0}}{v_{s}[1+\alpha_{1}]}\biggl\{1-\frac{\alpha_{1}(2\sigma r_{s})^{2}}{2[1+\alpha_{1}]}\biggr\},$ (75)
where the effective bubble radius $R=[\sqrt{1+\alpha_{1}}]/[2\sigma\sqrt{\alpha_{1}}]$ is defined by the criterion $f(R)\approx f(0)/2$, which describes the smooth decay of the spatial distortion. This consideration is analogous to that contemplated by Alcubierre [45], where $R$ is determined by the point at which $f(r_{s})$ reaches a significant fraction of its maximum value (not explicitly 1/2, but analogous). This condition, although arbitrary, marks the limit where the spacetime perturbation becomes negligible, ensuring a clear physical transition between the distorted and flat regions.
In the second case, we consider the region external to the bubble, $|r_{s}|\gg R\rightarrow\cosh{(2\sigma r_{s})}\approx\exp{(2\sigma r_{s})}/2$, which in this regime of small perturbations the function $f(r_{s})$ admits the following approximation
$f(r_{s})\approx\frac{2\alpha_{0}}{v_{s}\alpha_{1}}\exp{[-(2\sigma r_{s})]}.$ (76)
On the other hand, analyzing the quantum potential $Q_{II}$ (Eq. 48), we verify that it acts as an effective energy density ($Q_{II}]=\mbox{Energy}/\mbox{Length}$ in natural units), unlike classical Bohmian mechanics. Furthermore, in the context of the Alcubierre metric, the predominant term in Eq. (48) in this regime of small deformations is given by:
$Q_{II}\approx\frac{v_{s}^{2}}{2}f^{2}(r_{s}),$ (77)
where $f(r_{s})$ describes the bubble profile. This formulation implies that:
  • 1.
    $Q_{II}$ generates a spacetime distortion whose total energy,
    $\mathcal{E}\approx\frac{v_{s}^{2}}{2}\int_{-a/2}^{a/2}dxf^{2}(r_{s}),$ (78)
    scales with the barrier width $a$;
  • 2.
    For $a\gg R$, the increase in distortion volume raises $\mathcal{E}$, which in turn amplifies $v_{s}$;
  • 3.
    There is a geometric coupling that ensures $v_{s}\propto a$, keeping $\Delta t$ constant (Hartman effect).
Consequently, $Q_{II}$ not only guides trajectories, but becomes an active source of spacetime deformation. Then, for narrow barriers $a\leq R$, in this regime of small perturbations ($\sigma\ll 1$), we obtain
$\mathcal{E}_{\leq}$ $\approx$ $\frac{v_{s}^{2}}{2}\int_{-a/2}^{a/2}dx \frac{\alpha_{0}^{2}}{v_{s}^{2}[1+\alpha_{1}]^{2}}\biggl\{1-\frac{\alpha_{1}(\cancelto{\ll 1}{2\sigma r_{s}})^{2}}{2[1+\alpha_{1}]}\biggr\}^{2}$
$=$ $\frac{\alpha_{0}^{2}}{2[1+\alpha_{1}]^{2}}\int_{-a/2}^{a/2}dx$
$=$ $\frac{\alpha_{0}^{2}a}{2[1+\alpha_{1}]^{2}},$
which, upon substituting $\alpha_{0}\approx-v_{s}$ and $\alpha_{1}\approx 1$, reduces to
$\mathcal{E}_{\leq}\approx\frac{v_{s}^{2}a}{8}.$ (79)
Similarly, for wide barriers $a>2R$, in this regime of small perturbations, we obtain
$\mathcal{E}_{>}$ $\approx$ $\frac{v_{s}^{2}}{2}\biggl\{\frac{4\alpha_{0}^{2}}{v_{s}^{2}\alpha_{1}^{2}}\exp{(4\sigma x_{s})}\int_{-a/2}^{-R}dx\exp{(-4\sigma x)}$
$+$ $\int_{-R}^{R}dxf^{2}(0)+\frac{4\alpha_{0}^{2}}{v_{s}^{2}\alpha_{1}^{2}}\exp{(4\sigma x_{s})}\int_{R}^{a/2}dx\exp{(-4\sigma x)}\biggr\}$
$=$ $\frac{v_{s}^{2}}{2}\biggl\{\frac{4\alpha_{0}^{2}}{v_{s}^{2}\alpha_{1}^{2}}\exp{(4\sigma x_{s})}\int_{-a/2}^{-R}dx\exp{(-4\sigma x)}$
$+$ $\frac{\alpha_{0}^{2}}{v_{s}[1+\alpha_{1}]^{2}}\int_{-R}^{R}dx+\frac{4\alpha_{0}^{2}}{v_{s}^{2}\alpha_{1}^{2}}\exp{(4\sigma x_{s})}\int_{R}^{a/2}dr_{s}\exp{(-4\sigma x)}\biggr\}.$
Upon integration and some algebraic procedures, this equation reduces to
$\mathcal{E}_{>}$ $\approx$ $\alpha_{0}^{2}\biggl\{\frac{R}{[1+\alpha_{1}]^{2}}+\frac{\exp{(4\sigma x_{s})}\exp{(-4\sigma R)}}{\alpha_{1}^{2}\sigma}$ (81)
- $\frac{\exp{(4\sigma x_{s})}\exp{(-2\sigma a)}}{\alpha_{1}^{2}\sigma}\biggr\},$
from which, after considering $\sigma\ll 1\rightarrow 4\sigma R\approx 0$, $4\sigma x_{s}\approx 0$, $a\gg R\rightarrow 2\sigma a\gg 2\sigma R$ and substituting $\alpha_{0}\approx-v_{s}$ and $\alpha_{1}\approx 1$, we finally obtain
$\mathcal{E}_{>}\approx\frac{v_{s}^{2}}{\sigma}.$ (82)
Therefore, for narrow barriers, $a\leq R$, from Eq. (79) we obtain a velocity $v_{s}\approx\sqrt{8\mathcal{E}_{\leq}/a}$, which implies a tunneling time $\Delta t=3[a^{3}/8\mathcal{E}_{\leq}]^{1/2}$ (see figure 3), which clearly depends on $a$ (there is no Hartman effect). But for very wide barriers, $a>2R$, from Eq. (82) we obtain a velocity $v_{s}\approx\sqrt{\sigma\mathcal{E}_{>}}$, and since $v_{s}\propto a\rightarrow\sqrt{\sigma\mathcal{E}_{>}}\propto a$. Because of this, the energy must scale as $\mathcal{E}_{>}=[n_{0}a]^{2}/\sigma$ (where $n_{0}$ is a proportionality constant) so that the bubble velocity scales as $v_{s}=n_{0}a$ (see Fig. 5).
Thus, we obtain the constant tunneling time $\Delta t=3/n_{0}$ (see Fig. 4), in which it is evident that the energy-width coupling implies that as the barrier widens, the quantum potential provides exactly the additional distortion energy needed to maintain a fixed tunneling time. The Alcubierre-type spacetime, Eq. (33), mediates this process through the geometric relationship between $Q_{II}$ and $v_{s}$ in Eq. (48), achieving superluminal tunneling (even for low-energy incident particles $E\ll V_{0}$) without violating relativistic principles [40, 45]. This explains the saturation of tunneling times (Hartman effect) observed in Refs. [1, 17], through a geometric compensation mechanism.
Figure 3: Tunneling time ($\Delta t=3[a^{3}/8\mathcal{E}_{\leq}]^{1/2}$) for narrow barriers ($a\leq R$) with different energy values ($\mathcal{E}_{\leq}$).
Figure 4: Constant tunneling time ($\Delta t=3/n_{0}$) (Hartman effect) for wide barriers ($a>2R$) with different scale factors ($n_{0}$).
Figure 5: Linear relationship between $v_{s}/c$ and barrier width $a$. The blue area indicates the superluminal regime ($v_{s}/c>1$), reached when $a>c/n_{0}\approx 1.4R$.